# EMMA: Efficient Visual Alignment in Multi-Modal LLMs

### v1

@@ 3.2 Analysis on Modality Adaptation by EMMA
As explained earlier, the Visual Alignment module, which consists of a linear layer, forms the core of our proposed method. Let $f:\mathbb{R}^{(n+m)\times d}\rightarrow\mathbb{R}^{n\times d}$ represent the Visual Alignment module, where $n$ and $m$ denote the number of visual and textual tokens, respectively, each with a dimensionality of $d$. The Visual Alignment module takes the concatenation of the visual and textual tokens as input to generate $n$ refined visual tokens, and we begin our examination by scrutinizing the matrix $W\in\mathbb{R}^{(n+m)\times n}$ associated with it. By analyzing the norms of the weights corresponding to each token, we can identify which tokens, visual or textual, are most impactful. The histogram of $\ell_{2}$ and $\ell_{1}$ norms of visual and textual tokens are demonstrated in Figure 5. As the weights of textual tokens are below 1 and the weights of visual tokens are above 1, the $\ell_{1}$ norm of the weights serve as a more indicative measure of token importance. As expected, the visual tokens exhibit higher weights within the Visual Alignment module, signifying their greater influence on the generated multi-modal representation. Another key observation is that certain textual tokens exert more influence than others. To highlight the relative importance of each textual token, Figure 5(a) presents a bar chart illustrating their respective weights. CLIP’s text encoder generates 77 textual tokens, shown along the x-axis. As depicted in the figure, the earlier tokens tend to have a more significant impact. This finding indicates that the alignment module effectively identifies the most informative textual tokens, as instructions typically consist of brief prompts, with the crucial information concentrated in the early tokens and the remaining tokens often masked.
Figure 6: We evaluate the effectiveness of the Visual Alignment module by analyzing how well it refines visual representations to capture instruction-specific details. This is measured by calculating the mutual information between the visual representations and the response encodings produced by the text encoder. As illustrated in the figure, EMMA’s mutual information is 1.5 times higher than LLaVA’s, demonstrating its superior alignment with the instruction.
Enhancement in Modality Alignment.
The primary objective of EMMA’s modality adaptation is to align visual representations with the instructions, ensuring they emphasize the aspects of the image directed by the instructions. Our method achieves this by incorporating instruction encodings into the refinement process of the visual representations. In this section, we explore the alignment capabilities of EMMA by examining the visual representations before and after alignment. To perform this analysis, we utilize the MMVP Tong et al. (2024) benchmark, which is designed to expose the visual shortcomings of MLLMs. We focus on images with visually similar encodings but subtle differences, as shown in Figure 4. EMMA demonstrates a 9.3% improvement on this benchmark, underscoring its ability to generate more distinguishable visual representations for such images. To empirically validate this, we compute the $\ell_{2}$ norm of visual representations between randomly selected pairs of MMVP images, comparing the results for pre-aligned and post-aligned representations. The histogram of the norms, shown in Figure 5(b), reveals a clear shift, indicating that the aligned representations are better at distinguishing between these images by focusing on instruction-relevant tokens.
Mutual Information between Aligned Visual Tokens and Response Tokens.
@@
Another primary objective of the Modality Adaptation module is to adapt the visual representations with the language modality, ensuring they are well-aligned with the language model and encapsulate the necessary information to accurately respond to the given prompt. To evaluate the contribution of visual tokens to the language model’s output, we use the Mutual Information which quantifies the amount of information obtained about one random variable through the other. In this analysis, the LLaVA-In-Wild Liu et al. (2024b) benchmark is employed, which has a set of 24 images with 60 challenging questions in novel domains. For each of the 60 samples, visual representations are generated using the visual modules of both LLaVA and EMMA. It is important to note that EMMA’s visual module processes the image and prompt to produce instruction-aware representations, whereas LLaVA generates instruction-agnostic encodings. Additionally, for each sample, the corresponding answer is encoded using the Text Encoder. The Mutual Information between the visual and response encodings is then calculated, with the results shown in Figure 6. As illustrated, the mean mutual information for EMMA is 1.5 times higher than that of LLaVA, underscoring the effectiveness of EMMA’s Visual Alignment in steering the model toward accurate language responses.
[…]
BLIP-2 Vicuna-13B 13B 129M 65.0 19.6 61 41 45.9
InstructBLIP Vicuna-7B 7.91B 130.2M - 34.5 60.5 49.2 -
InstructBLIP Vicuna-13B 14.2B 130.2M - 33.4 63.1 49.5 -
Shikra Vicuna-13B 13b 6.1M 77.4 - - - 47.2
IDEFICS LLaMA-7B 9B 354M 50.9 35.5 - 38.4 -
IDEFICS LLaMA-65B 80B 354M 60.0 36.0 - 45.2 -
Qwen-VL-Chat Qwen-7B 9.6B 1.4B 78.2 38.9 68.2 57 56.6
LLaVA-1.5 Vicuna-7B 7B 1.2M 78.5 50.0 66.8 62.0 -
mPLUG-Owl2 LLaMA-7B 8.2B 348M 79.4 54.5 68.7 56.1 57.7
BRAVE FlanT5-XL 10B 100M 82.5 54.2 - 52.7 66.0
EMMA Vicuna-7B 7B 1.8M 89.42 56.03 73.14 56.01 68.57
Table 2: Comparison between EMMA and previous methods on academic task-oriented datasets. EMMA achieves state-of-the-art performance on 4/5 of the benchmarks while using the fewest number of parameters and training data compared to the previous approaches, utilizing modality adaptation. (PT and IT stand for Pretraining and Instruction fine-tuning repectively.)
@@ 4 Experimental Evaluation
We employ CLIP-ViT-L-14, trained on 3362 pixel images, as the base image encoder and text encoder. The Visual Alignment module is initialized with the identity matrix for the visual tokens and all zeros for the instruction tokens to transfer all the visual tokens at the beginning of training. Moreover, the Visual Alignment module is designed to maintain the same number of visual tokens as the baseline model. The latest Vicuna v1.5 Zheng et al. (2023) is used as the base LLM. EMMA uses the same set of hyperparameters as the LLaVA-1.5. For all the analysis performed in the Section 3, we use the same dataset as the baseline model, which is 558K and 665K samples for the pretraining and fine-tuning stages respectively. In the Evaluation setting, we have preserved the same pertaining data but scaled the fine-tuning data to 1.2M samples, which is the most efficient data compared to all of the state-of-the-art methods (except for LLaVA) as shown in Table 4.
Benchmarks.
@@
We evaluate EMMA across 10 benchmarks that span a diverse set of tasks, including scientific question answering, visual question answering with poor image quality, integrated perception and reasoning tasks, visual dialogue, and general reasoning. The results of recent benchmarks designed specifically for instruction-following large multi-modal models (LMMs) are presented in Table 4, while Table 2 outlines the performance on academic task-oriented benchmarks. EMMA emerges as the state-of-the-art model for 4 out of 5 academic task-oriented benchmarks, outperforming mPLUG-Owl2 which has 50×50$\times$50 × larger modality adaptation module and is trained on 300×300$\times$300 × more data, and BRAVE which has 24×24$\times$24 × larger vision encoder and is trained on 100×100$\times$100 × more data. On MLLM specialized benchmarks, EMMA delivers the best performance on 3 out of 5 benchmarks and second best on the two others with less than 0.5% difference. In conclusion, EMMA achieves the best performance in 7 out of 10 benchmarks, despite utilizing the simplest architecture, outperforming other MLLMs that rely on complex modality adaptation modules.
Model AMBER FOIL
Attr Rel Exis
LLaVA-1.5 20.23 37.1 91.43 50.08
EMMA-1.0 26.15$+$5.92 37.34$+$0.24 94.50$+$3.07 66.13$+$16.05
Table 3: Robustness to Hallucination. This table compares EMMA’s robustness against hallucinations with LLaVA, demonstrating consistent improvements over the baseline in avoiding hallucinations.
@@
Robustness and the ability to avoid hallucination are essential for multi-modal large language models (MLLMs), which are increasingly applied in areas like medical diagnostics to interpret complex text and image data. Hallucination, a significant security threat to MLLMs, occurs when the model generates information that does not accurately represent the provided images or text. As a result, evaluating and mitigating hallucinations is a critical step in ensuring the reliability of MLLMs before real-world deployment, making it a key focus in model performance assessments. The hallucination evaluations in this section are performed utilizing two benchmarks, AMBER and FOIL, which do not rely on additional LLMs, thereby providing a direct and controlled means of assessing the model’s ability to avoid hallucinations. These benchmarks focus on specific challenges in multi-modal reasoning, testing the model’s accuracy in aligning textual and visual content without introducing erroneous information.
Figure 7: This radar chart compares EMMA’s performance when leveraging textual features generated by either the penultimate or last layer of the text encoder, highlighting the advantage of features associated with the penultimate layer.
AMBER consists of 7628, 4924, and 1663 samples for the Attribute, Relation, and Existence categories, respectively. FOIL contains a total of 99,480 test samples, of which 92,705 are straightforward, allowing both LLaVA and our method to successfully avoid hallucinations. However, 6,775 samples present more challenging cases. We compare LLaVA-1.5, the baseline for our approach, with EMMA. The results, shown in Table 3, indicate that EMMA outperforms the baseline, with significant performance gaps in two of the four benchmarks.
@@
We ablate the text feature’s abstraction level in this section. The textual features can be optionally derived from either the final layer or the penultimate layer of the CLIP Text Encoder. A comparison between the two methods of extracting textual features is presented in Figure 7, demonstrating the clear superiority of features derived from the penultimate layer. We hypothesize that the final layer of CLIP captures more global and abstract semantics of the instruction, whereas the penultimate layer focuses on finer details. Additionally, since visual features are extracted from the corresponding layer in the visual encoder, using the same layer in the text encoder ensures both modalities operate at a similar level of abstraction, promoting better alignment between them.
@@ 5 Conclusion
In this work, we addressed the inefficiencies in the modality adaptation modules employed by current multi-modal large language models. We hypothesized that the initial alignment between visual and textual encodings plays a critical role in determining the complexity level of the modality adaptation module and the necessary amount of training data. Our lightweight approach, EMMA (Enhanced Multi-Modal Adaptation), leverages CLIP’s text encoder to generate instruction encodings and, by exploiting this initial alignment, demonstrates that the modality adaptation module can be simple while still enhancing the alignment between visual and textual modalities. Through extensive analysis, we demonstrated that EMMA effectively produces instruction-aware visual representations aligned with the language model. Our experiments, evaluated across multiple benchmarks, show that EMMA significantly outperforms state-of-the-art models that use modality adaptation modules 50×50$\times$50 × larger. Finally, our robustness analysis, particularly in hallucination avoidance, confirmed EMMA’s superior ability to accurately process multi-modal data, even in the challenging scenarios.

### v2

@@ 3.2 Analysis on Modality Adaptation by EMMA
The Mechanics of the Visual Alignment Module. The Visual Alignment module takes the concatenation of the visual and textual tokens ($n+m$ tokens) as input to generate the $n$ refined visual tokens. We begin our examination by scrutinizing the matrix $W\in\mathbb{R}^{(n+m)\times n}$ associated with it. By analyzing the norms of the weights corresponding to each token, we can identify which tokens, visual or textual, are most impactful. The histogram of $\ell_{1}$ norm of visual and textual tokens are demonstrated in Figure 2. As the weights of textual tokens are below 1 and the weights of visual tokens are above 1, the $\ell_{1}$ norm of the weights serve as a more indicative measure of token importance. As expected, the visual tokens exhibit higher weights within the Visual Alignment module, signifying their greater influence on the generated multi-modal representation. Another key observation is that certain textual tokens exert more influence than others. To highlight the relative importance of each textual token, Figure 2(b) presents a bar chart illustrating their respective weights. CLIP’s text encoder generates 77 textual tokens, shown along the x-axis. As depicted in the figure, the earlier tokens tend to have a more significant impact. This finding indicates that the alignment module effectively identifies the most informative textual tokens, as instructions typically consist of brief prompts, with the crucial information concentrated in the early tokens and the remaining tokens often masked.
@@
Mutual Information between Aligned Visual Tokens and Response Tokens. Another primary objective of the Modality Adaptation module is to adapt the visual representations to the language modality, ensuring they are well-aligned with the language model and encapsulate the necessary information to accurately respond to the given prompt. To evaluate the contribution of visual tokens to the language model’s output, we use the Mutual Information, which quantifies the amount of information obtained about one random variable through the other. In this analysis, the LLaVA-In-Wild Liu et al. (2024b) benchmark is employed, which has a set of 24 images with 60 challenging questions in novel domains. For each of the 60 samples, visual representations are generated using the visual modules of both LLaVA and EMMA. It is important to note that EMMA’s visual module processes the image and prompt to produce instruction-aware representations, whereas LLaVA generates instruction-agnostic encodings. Additionally, for each sample, the corresponding answer is encoded using CLIP’s text encoder. The Mutual Information between the visual and response encodings is then calculated, with the results shown in Figure 3(a). As illustrated, the mean mutual information for EMMA is 1.5 times higher than that of LLaVA, underscoring the effectiveness of EMMA’s Visual Alignment in steering the model toward accurate language responses.
Isolated Impact of the Modality Adaptation Module.
To evaluate the improvements introduced by EMMA’s modality adaptation, we follow the same training scheme as the baseline model (Liu et al., 2024a). Therefore the improvements demonstrated in Figure 3(b) are achieved solely by our proposed method, EMMA. Across multiple benchmarks compared to the baseline model, LLaVA-1.5. EMMA consistently surpasses LLaVA-1.5 in various cross-modal tasks, demonstrating significant improvements in MMVP (Tong et al., 2024) (+9.3%), MuirBench (Wang et al., 2024a) (+5.62%), SQA (Lu et al., 2022b) (+4.2%), MMBench (Liu et al., 2023a) (+1.9%), and POPE (Li et al., 2023e) (+1.17%). These results underscore EMMA’s effectiveness in enhancing cross-modal tasks that require both visual and textual comprehension.
(a) Instances of MMVP Benchmark
(b) $\ell_{2}$ distance between the MMVP image pairs.
Aligned vs. Vanilla Visual Tokens The primary objective of EMMA’s modality adaptation is to align visual representations with the instructions, ensuring they emphasize the aspects of the image directed by the instructions. Our method achieves this by incorporating instruction encodings into the refinement process of the visual representations. In this section, we explore the alignment capabilities of EMMA by examining the visual representations before and after alignment. To perform this analysis, we utilize the MMVP Tong et al. (2024) benchmark, which is designed to expose the visual shortcomings of MLLMs. We focus on images with visually similar encodings but subtle differences, as shown in Figure 4(a). EMMA demonstrates a 9.3% improvement on this benchmark, underscoring its ability to generate more distinguishable visual representations for such images. To empirically validate this, we compute the $\ell_{2}$ norm of visual representations between randomly selected pairs of MMVP images, comparing the results for pre-aligned and post-aligned representations. The histogram of the norms, shown in Figure 4(b), reveals a clear shift, indicating that the aligned representations are better at distinguishing between these images by focusing on instruction-relevant tokens.
@@ 4 Experimental Evaluation
For training, we follow the same two-stage instruction fine-tuning process as LLaVA. In the pretraining stage, only the Visual Alignment and Projection modules are trained, while the language model remains frozen. During the fine-tuning stage, the LLM is unfrozen and fine-tuned along with the two aforementioned modules. We employ CLIP-ViT-L-14, trained on 3362 pixel images, as the base image encoder and text encoder. The Visual Alignment module is initialized with the identity matrix for the visual tokens and all zeros for the instruction tokens to transfer all the visual tokens at the beginning of training. Moreover, the Visual Alignment module is designed to maintain the same number of visual tokens as the baseline model. The latest Vicuna v1.5 (Zheng et al., 2023) is used as the base LLM. EMMA uses the same set of hyperparameters as the LLaVA-1.5. For all the analysis performed in the Section 3, we use the same dataset as the baseline model, which is 558K and 665K samples for the pretraining and fine-tuning stages, respectively. In the Evaluation setting, we have preserved the same pertaining data but scaled the fine-tuning data to 1.2M samples, including LVIS-Instruct4V(Wang et al., 2023a), CLEVR(Johnson et al., 2017), VizWiz(Gurari et al., 2018), and ScienceQA(Lu et al., 2022a) training data. EMMA’s training data is the most efficient compared to all state-of-the-art methods (except for LLaVA), as shown in Table 3.
[…]
BLIP-2 Vicuna-13B ViT-g/14 13B 129M 65.0 19.6 61 41 45.9
InstructBLIP Vicuna-7B ViT-g/14 7.91B 130.2M - 34.5 60.5 49.2 -
InstructBLIP Vicuna-13B ViT-g/14 14.2B 130.2M - 33.4 63.1 49.5 -
Shikra Vicuna-13B ViT-L/14 13b 6.1M 77.4 - - - 47.2
IDEFICS LLaMA-7B ViT-H/14 9B 354M 50.9 35.5 - 38.4 -
IDEFICS LLaMA-65B ViT-H/14 80B 354M 60.0 36.0 - 45.2 -
Qwen-VL-Chat Qwen-7B ViT-G/14 9.6B 1.4B 78.2 38.9 68.2 57 56.6
LLaVA-1.5 Vicuna-7B ViT-L/14 7B 1.2M 78.5 50.0 66.8 62.0 -
QA-ViT Vicuna-7B ViT-L/14 7B 1.2M 80.5 36.5 - - -
mPLUG-Owl2 LLaMA-7B ViT-L/14 8.2B 348M 79.4 54.5 68.7 56.1 57.7
BRAVE FlanT5-XL $\clubsuit$ 10B 100M 82.5 54.2 - 52.7 66.0
EMMA Vicuna-7B ViT-L/14 7B 1.8M 89.42 56.03 73.14 56.01 68.57
Table 2: Comparison between EMMA and previous methods on academic task-oriented datasets. EMMA achieves state-of-the-art performance on 4/5 of the benchmarks while using the fewest number of parameters and training data compared to the previous approaches, utilizing modality adaptation(PT and IT stand for Pretraining and Instruction fine-tuning, respectively). The $\clubsuit$ symbol denotes the collection of vision encoders utilized by BRAVE, including EVA-CLIP-g Fang et al. (2023), ViT-L/14 Radford et al. (2021), SILC-G/16 Naeem et al. (2025), ViT-e Chen et al. (2022), and DINOv2-L/14 Oquab et al. (2023).
@@
Benchmarks. We evaluate EMMA across 10 benchmarks that span a diverse set of tasks, including scientific question answering, visual question answering with poor image quality, integrated perception and reasoning tasks, visual dialogue, and general reasoning. The results of recent benchmarks designed specifically for instruction-following large multi-modal models (LMMs) are presented in Table 3, while Table 2 outlines the performance on academic task-oriented benchmarks. EMMA emerges as the state-of-the-art model for 4 out of 5 academic task-oriented benchmarks, outperforming mPLUG-Owl2 which has 50×50$\times$50 × larger modality adaptation module and is trained on 300×300$\times$300 × more data, and BRAVE which has 24×24$\times$24 × larger vision encoder and is trained on 100×100$\times$100 × more data. On MLLM specialized benchmarks, EMMA delivers the best performance on 3 out of 5 benchmarks and is second best on the two others with less than 0.5% difference. In conclusion, EMMA achieves the best performance in 7 out of 10 benchmarks, despite utilizing the simplest architecture, outperforming other MLLMs that rely on complex modality adaptation modules.
@@
Robustness and the ability to avoid hallucination are essential for multi-modal large language models (MLLMs), which are increasingly applied in areas like medical diagnostics to interpret complex text and image data. Hallucination, a significant security threat to MLLMs, occurs when the model generates information that does not accurately represent the provided images or text. As a result, evaluating and mitigating hallucinations is a critical step in ensuring the reliability of MLLMs before real-world deployment, making it a key focus in model performance assessments. The hallucination evaluations in this section are performed utilizing two benchmarks, AMBER and FOIL, which do not rely on additional LLMs, thereby providing a direct and controlled means of assessing the model’s ability to avoid hallucinations. These benchmarks (Wang et al., 2023b; Shekhar et al., 2017; Nie et al., 2024; Wu et al., 2024) focus on specific challenges in multi-modal reasoning, testing the model’s accuracy in aligning textual and visual content without introducing erroneous information. AMBER consists of 7628, 4924, and 1663 samples for the Attribute, Relation, and Existence categories, respectively. FOIL contains a total of 99,480 test samples, of which 92,705 are straightforward, allowing both LLaVA and our method to successfully avoid hallucinations. However, 6,775 samples present more challenging cases. We compare LLaVA-1.5, the baseline for our approach, with EMMA. The results, shown in Figure 5(a), indicate that EMMA outperforms the baseline, with significant performance gaps in two of the four benchmarks.
[…]
BLIP-2 Vicuna-13B ViT-g/14 14.2B 129M - - - - -
InstructBLIP Vicuna-7B ViT-g/14 7.91B 130.2M 36 23.7 - - -
InstructBLIP Vicuna-13B ViT-g/14 14.2B 130.2M - - - 25.3 -
Shikra Vicuna-13B ViT-L/14 13B 6.1M 58.8 - - - -
IDEFICS LLaMA-7B ViT-H/14 9B 354M 48.2 25.2 - 19.8 -
IDEFICS LLaMA-65B ViT-H/14 80B 354M 54.5 38.1 - - -
Qwen-VL-Chat Qwen-7B ViT-G/14 9.6B 1.4B 60.6 56.7 35.9 - -
LLaVA-1.5 Vicuna-7B ViT-L/14 7B 1.2M 64.3 58.3 35.11 21.1 23.46
mPLUG-Owl2 LLaMA-7B ViT-L/14 8.2B 348M 64.5 - 32.7 22.2 -
BRAVE FlanT5-XL $\clubsuit$ 10B 100M - - - -
EMMA Vicuna-7B ViT-L/14 7B 1.8M 66.44 60.15 35.44 25.1 32.0
Table 3: This table compares EMMA with previous methods on specialized MLLM benchmarks. EMMA delivers the best performance on 3/5 benchmarks and secures second place on the remaining two with less than 0.5% difference. The $\clubsuit$ symbol denotes the collection of vision encoders utilized by BRAVE, including EVA-CLIP-g Fang et al. (2023), ViT-L/14 Radford et al. (2021), SILC-G/16 Naeem et al. (2025), ViT-e Chen et al. (2022), and DINOv2-L/14 Oquab et al. (2023).
@@
We ablate the text feature’s abstraction level in this section. The textual features can be optionally derived from either the final layer or the penultimate layer of the CLIP Text Encoder. A comparison between the two methods of extracting textual features is presented in Figure 5(b), demonstrating the clear superiority of features derived from the penultimate layer. We hypothesize that the final layer of CLIP captures more global and abstract semantics of the instruction, whereas the penultimate layer focuses on finer details. Additionally, since visual features are extracted from the corresponding layer in the visual encoder, using the same layer in the text encoder ensures both modalities operate at a similar level of abstraction, promoting better alignment between them.
Figure 5: Left: This plot compares EMMA’s robustness against hallucinations with LLaVA, demonstrating consistent improvements over the baseline in avoiding hallucinations. Right: This radar chart compares EMMA’s performance when leveraging textual features generated by either the penultimate or last layer of the text encoder, highlighting the advantage of features associated with the penultimate layer.
Ablation on the Visual Alignment Module.
To evaluate the effectiveness of our proposed visual alignment module, implemented as a linear layer, we conduct a comparative analysis by replacing the linear layer with a cross-attention mechanism between the two modalities. Specifically, we compare the performance of three MLLMs, all sharing the same architecture and training setup but differing in their modality adaptation modules:
  • LLaVA – a baseline model without any modality adaptation,
  • EMMA (Linear) – utilizing a linear layer for modality adaptation,
  • EMMA (Cross-Attention) – using cross-attention for modality adaptation.
The results in Table 4 show that EMMA with the linear adaptation module consistently achieves the best performance across most benchmarks. Moreover, EMMA with cross-attention shows improvements over the baseline on SQA and MuirB, highlighting the potential of its adaptation mechanism. Nevertheless, its more complex architecture introduces risks of overfitting, as observed in benchmarks such as GCA, MMB, MMVP, and A-Bench. We posit that effective modality adaptation, i.e., refining visual representations through instruction-based guidance, requires a carefully balanced interplay between architectural complexity and training set size. This balance becomes particularly crucial when operating under constraints such as our baseline setup with only 1.2 million training samples.
Method Data SQA GQA POPE MMB-en MMB-cn MMVP A-Bench MuirB
LLaVA-1.5 1.2M 66.80 62.00 86.10 64.30 58.30 24.70 62.97 23.46
EMMA (Linear) 1.2M 71.00 62.78 87.27 66.19 60.20 34.00 63.48 29.08
EMMA (Cross-Attention) 1.2M 67.08 49.33 79.73 60.84 51.42 8.67 56.01 29.50
Table 4: A comparative analysis of EMMA’s Visual Alignment module reveals that the linear adaptation layer delivers the most consistent performance gains. In contrast, the cross-attention variant demonstrates selective improvements but is prone to potential overfitting in certain scenarios.
@@ 5 Conclusion
In this work, we addressed the inefficiencies in the modality adaptation modules employed by current multi-modal large language models. We hypothesized that the initial alignment between visual and textual encodings plays a critical role in determining the complexity level of the modality adaptation module and the necessary amount of training data. Our lightweight approach, EMMA (Enhanced Multi-Modal Adaptation), leverages CLIP’s text encoder to generate instruction encodings and, by exploiting this initial alignment, demonstrates that the modality adaptation module can be simple while still enhancing the alignment between visual and textual modalities. Through extensive analysis, we demonstrated that EMMA effectively produces instruction-aware visual representations aligned with the language model. Our experiments, evaluated across multiple benchmarks, show that EMMA significantly outperforms state-of-the-art models that use modality adaptation modules 50×50$\times$50 × larger. Finally, our robustness analysis, particularly in hallucination avoidance, confirmed EMMA’s superior ability to accurately process multi-modal data, even in challenging scenarios.